OpenStudy (anonymous):

Which of the following is equal to (2x/3 - 7)+7? (2x-7)-21 2x-21/3 3x/2 2x/3

OpenStudy (anonymous):

Hi: \[(\left(\begin{matrix}2 \\3\end{matrix}\right)x-7)+7=2\frac{ x }{ 3}\]

OpenStudy (anonymous):

Thanks so much!